# SCott: Accelerating Diffusion Models with Stochastic Consistency Distillation

### v1

@@ 5.1 Comparison with Other Methods
Table 3: Comparisons with the state-of-the-art methods on MJHQ-5K in terms of FID, CS, CR. All models are based on RV5.1.
Method Step Time (s) FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
DPM++ (Lu et al., 2022b) 25 0.88 22.1 0.320 0.8664
DDIM (Song et al., 2020b) 50 -- 22.9 0.320 0.8536
LCM (Luo et al., 2023a) 2 0.13 37.2 0.296 0.7016
Ours 2 0.13 24.9 0.301 0.8794
Comparison on MJHQ-5K with RV5.1. To better assess the quality of produced images, we further conduct experiments on MJHQ-5K, randomly selected from MJHQ-30K [^3] https://huggingface.co/playgroundai/playground-v2-1024px-aesthetic., since it owns high image quality and high image-text alignment, and the correlation between human preference and FID score on the MJHQ is verified by user study. The evaluation metrics of several methods with RV5.1 as a teacher are listed in Table 3. The reason for distilling RV5.1 is that RV5.1 is much stronger than SD1.5 for text-to-image consistency, which can be found in Table 8. LCM is also implemented according to the official code. Since the training codes of PD, CAD, InstaFlow, and UFOGen are unavailable, the metrics of these methods are not included. The result again presents our SCott outperforms LCM (Luo et al., 2023a) by a large margin, because our SDE-based CD provides a stronger and more versatile teacher than ODE-based CD in LCM, and the student output is further refined to be real data by the proposed discriminator.
5.2 Ablation Study
SDE solver. Table 4 depicts the results using deterministic and stochastic solvers for CD. DDIM($\eta$) represents DDIM solver with noise coefficient $\eta$, where $\eta$ denotes the hyper-parameter that controls the strength of random noise injected in Eq.2. $\eta$ achieves an interpolation between the deterministic DDIM ($\eta=0$) and original DDPM ($\eta$ = 1). As observed in Table 4, the SDE-based DDIM ($\eta=0.1,0.2$) surpasses ODE-based DDIM ($\eta=0$) for CD, demonstrating the superiority of SDE solver for CD. This is because the injection of noise at moderate intensity in SDE solvers aids in correcting estimated errors of the teacher model with multiple sampling steps, leading to a more powerful teacher. However, excessive noise intensity leads to poor convergence and degraded samples, as shown in DDIM($\eta=0.6$). These results indicate it is crucial to control the noise strength in SCott since large noise leads to low training stability. The 2-order ER-SDE 5 solver further enhances the performance, this is because 1. the 2-order ER-SDE 5 exhibits smaller discretization errors compared to the 1-order DDIM solver. 2. the noise function in ER-SDE 5 mitigates excessive noise, which leads to good convergence.
Table 4: The performance comparison of different solvers on MJHQ-5K with 2-step inference. All models are initialized by RV5.1.
@@ 5.2 Ablation Study
Table 5: The performance comparison of SDE solver steps with 2-step inference.
@@
Figure 4: Ablation study on adversarial training in SCott with varying inference steps.
Discriminator. Figure 4 illustrates our discriminator leads to gains with respect to FID and CS using different inference steps. The gains come from that incorporating adversarial training into SCott, our method introduces an error correction signal from the proposed discriminator to refine the outputs of student ${\bm{f}}_{\bm{\theta}}$, guiding the accelerated model ${\bm{f}}_{\bm{\theta}}$ to yield images close to the real data.
Inference steps.
Table 6: Performance comparison of our Scott on MJHQ-5K with different sampling steps.
Step FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
1 29.0 0.290 0.8226
2 24.9 0.301 0.8794
3 24.8 0.303 0.8804
4 24.7 0.304 0.8838
Table 6 demonstrates SCott consistently improves sample quality as NFE increases, including FID, CS, and CR, which benefit from the capability of CMs. Therefore, SCott achieves a trade-off between computational resources and quality. Visual results are shown in Section D.5.
6 Conclusions
@@
In this paper, we propose stochastic consistency distillation (SCott), a novel approach for accelerating text-to-image diffusion models. SCott integrates SDE solvers into consistency distillation to unleash the potential of the teacher, implemented by controlling noise strength and sampling step of SDE solvers. Adversarial learning is further utilized to aid SCott in generating high-quality images in rare-step sampling. SCott is capable of yielding high-quality images with 2 steps only, surpassing 2-step LCM and 1-step InstaFlow, comparable to 4-step UFOGen. Additionally, SCott consistently improves performance with increasing steps and exhibits higher diversity than competing baselines.
Impact Statement
There are many potential societal consequences of our work. Like other generative models, DPMs can be employed to generate misleading or harmful content, and our proposed method may further amplify the negative impact of generative AI for malicious purposes.
Appendix A Multi-step Sampling of Stochastic Consistency Distillation Model

### v4

@@ Introduction
Table 3: Comparisons with the state-of-the-art methods on MSCOCO-2014 30K in terms of FID with backbone SD1.5.
Comparison on MSCOCO-2017 5K with SD1.5
To kick-start the comparisons with the state-of-the-art methods, we evaluate the MSCOCO-2017 5K validation dataset (Lin et al. 2014). Zero-shot FID and CLIP Score (CS) with ViT-g/14 backbone (Radford et al. 2021) are exploited as objective metrics. To measure diversity, Coverage (CR) (Naeem et al. 2020) is used as another metric. Table 2 summarizes the performance of our SCott and comparative methods consisting of PD (Salimans and Ho 2022), CAD (Meng et al. 2023), LCM (Luo et al. 2023a), InstaFlow (Liu et al. 2023), UFOGen (Xu et al. 2024). Since PD (Salimans and Ho 2022), CAD (Meng et al. 2023), InstaFlow (Liu et al. 2023), and UFOGen (Xu et al. 2024) do not list CR value, we leave them --. Time denotes inference time on a single A100. LCM is implemented according to the official in our setting.
Method Step FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
DPM++ 25 22.1 0.320 0.8664
DDIM 50 22.9 0.320 0.8536
LCM 2 37.2 0.296 0.7016
SCott (Ours) 2 24.9 0.301 0.8794
Table 4: Comparisons with the state-of-the-art methods on MJHQ-5K in terms of FID, CS, CR, with backbone RV5.1.
Among all the methods, our 2-step SCott presents superior FID and CS values than 4-step PD, 8-step CAD, 1-step InstaFlow-0.9B, 2-step LCM. 2-step SCott achieves better FID than 1-step UFOGen and comparable CS. It is impressive to see that our 2-step SCott beats InstaFLow-1.7B which doubles the parameter size. These results demonstrate our SCott significantly enhances the quality of the generated images while reducing inference steps. The improvements lie in that the proposed $\mathcal{L_{CD}}$ in SCott is able to reduce the sampling steps remarkably and adversarial loss further improves image quality. Interestingly, the CS decreases by 0.004 when the inference step increases from 1 to 4 for UFOGen. Meanwhile, by changing the inference step from 1 to 4, all the metrics are improved for our SCott, and 4-step SCott outperforms 4-step UFOGen regarding both FID and CS. The result indicates the improvement is limited for UFOGen when increasing steps while our method is much more powerful in enhancing image quality with additional steps in a flexible manner, which is meaningful in the scenarios requiring high-quality images with more affordable computational budgets. The proposed SCott obtains higher CR than LCM, and even outperforms the ODE solvers, 25-step DPM++, and 50-step DDIM, indicating our SCott successfully increases sample diversity due to the introduced randomness in SDE solvers.
We present the qualitative comparisons involving InstaFlow and LCM (See Figure 5 in Appendix 5). Notably, 2-step SCott gains significant improvements over 2-step LCM and 1-step InstaFlow in terms of image quality and text-to-image alignment. We also observe that our generated images exhibit sharper textures and finer details, compared to the images generated by InstaFlow and LCM.
Comparison on MSCOCO-2014 30K with SD1.5 For complete comparisons, we also benchmark our method against DMD, SwiftBrush, and UFOGen on MSCOCO-2014 30K (Lin et al. 2014). We follow the evaluation setup in DMD and use a CFG scale of 3 following DMD. As depicted in Table 3, our 2-step SCott surpasses all the above methods in FID. Additionally, it’s worth noting that these methods can not improve output with more steps, whereas SCott offers the feasibility to enhance the sample quality with additional steps. With 4-step inference, SCott achieves a superior FID of 10.68. These results validate the effectiveness of our approach in achieving higher-quality samples within 2 steps and improving performance by leveraging additional steps.
@@
Table 5: Comparisons of different solvers in CD on MJHQ-5K with 2-step inference. All models are based on RV5.1.
Comparison on MJHQ-5K with RV5.1
To better assess the quality of produced images, we further conduct experiments on MJHQ-5K, randomly selected from MJHQ-30K [^3] https://huggingface.co/playgroundai/playground-v2-1024px-aesthetic., since it owns high image quality and image-text alignment, and the correlation between human preference and FID score on the MJHQ is verified by user study. The evaluation metrics of several methods with RV5.1 as a teacher are listed in Table 4. The reason for distilling RV5.1 is that RV5.1 is much stronger than SD1.5 for text-to-image consistency, which can be found in Appendix 5. LCM is also implemented according to the official code. Since the training codes of PD, CAD, InstaFlow, and UFOGen are unavailable, their metrics are not included. The result again presents our SCott outperforms LCM by a large margin, because our SDE-based CD provides a stronger and more versatile teacher than ODE-based CD in LCM, and the student output is further refined to be real data by the proposed discriminator.
Ablation Studies
SDE Solver. Table 5 depicts the results using deterministic and stochastic solvers for CD. DDIM ($\eta$) represents DDIM solver with noise coefficient $\eta$, where $\eta$ denotes the hyper-parameter that controls the strength of random noise injected in Eq.2. $\eta$ achieves an interpolation between the deterministic DDIM ($\eta=0$) and original DDPM ($\eta$ = 1). As observed in Table 5, the SDE-based DDIM ($\eta=0.1,0.2$) surpasses ODE-based DDIM ($\eta=0$) for CD, demonstrating the superiority of SDE solver for CD. This is because the injection of noise at moderate intensity in SDE solvers aids in correcting estimated errors of the teacher model with multiple sampling steps, leading to a more powerful teacher. However, excessive noise intensity leads to poor convergence and degraded samples, as shown in DDIM ($\eta=0.6$). These results indicate it is crucial to control the noise strength in SCott since large noise leads to low training stability. The 2-order ER-SDE 5 solver further enhances the performance, this is because 1. the 2-order ER-SDE 5 exhibits smaller discretization errors compared to the 1-order DDIM solver. 2. the noise function in ER-SDE 5 mitigates excessive noise, which leads to good convergence. Table 7 shows that SCott without GAN also surpasses the ODE solver-based LCM, further illustrating the benefits of using SDE solvers for CD.
@@
Table 6: Comparisons of choosing different SDE solver steps during training. All models are based on RV5.1 with 2-step inference on MJHQ-5K.
@@
Loss FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
SCott (Without GAN) 25.0 0.300 0.867
LCM 30.4 0.293 0.788
LCM + GAN 29.2 0.297 0.8160
SCott (Full Discriminator) 23.5 0.302 0.8880
SCott (Without Condition) 26.1 0.297 0.8412
SCott 21.9 0.310 0.9145
Table 7: The performance comparisons of Discriminator on MSCOCO-2017 5k with backbone SD 1.5.
Discriminator. Table 7 illustrates our discriminator leads to gains with respect to FID, CS, and CR using different inference steps. We observe our LoRA discriminator achieves better sample qualities than the fully parameterized one. Compared with the discriminator without condition, SCott leverages both the time and text conditions. Such a design helps the discriminator better distinguish between real and generated samples. Interestingly, we observe that our discriminator can also improve the outputs of LCM, which indicates the generalization capability of our proposed discriminator.
Conclusions
@@
In this paper, we propose stochastic consistency distillation (SCott), a novel approach for accelerating text-to-image diffusion models. SCott integrates SDE solvers into consistency distillation to unleash the potential of the teacher, implemented by controlling noise strength and sampling step of SDE solvers. Adversarial learning is further utilized to aid SCott in generating high-quality images in rare-step sampling. SCott is capable of yielding high-quality images with 2 steps only, surpassing 2-step LCM, 1-step InstaFlow, and 4-step UFOGen. Additionally, SCott consistently improves performance by increased inference cost within 4 steps and exhibits higher diversity than competing baselines.
Supplementary Materials
SCott: Accelerating Diffusion Models with Stochastic Consistency Distillation
Supplementary Materials
1 Theoretical justification for using SDE solvers in CD
Theorem 2.
Let $\Delta t:=\max\limits_{n\in[1,N]}{|t_{n+1}-t_{n}|}$ where $t\in[\tau,T]$. Assume ${\bm{f}}_{\theta}(\cdot,\cdot)$ is Lipschitz in ${\bm{x}}$ with constant $L_{1}$. Denote ${\bm{f}}(\cdot,\cdot)$ the consistency function of the reverse SDE defined in Equation 3 in the main part of our paper. Assume the SDE solver $\Phi_{SDE}$ has a local error bound of $O((t_{n+1}-t_{n}))^{p+1}$ with $p\geq 1$. Then, if we have $\mathcal{L}_{\mathcal{KL}}^{N}({\bm{\theta}},\Phi_{SDE})=0$, we have
$\sup\limits_{n,{\bm{x}}}||{\bm{f}}_{\theta}({\bm{x}}_{t_{n}},t_{n})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n})||_{2}=O((\Delta t)^{p})$
The following lemma provides the convergence proof of using ODE solvers in CD (proof in (Song et al. 2023)), which is crucial to our proof for Theorem 2.
Lemma 3 (Proof in (Song et al. 2023) ).
Let $\Delta t:=\max\limits_{n\in[1,N]}{|t_{n+1}-t_{n}|}$ where $t\in[\tau,T]$. Assume ${\bm{f}}_{\theta}(\cdot,\cdot)$ is Lipschitz in ${\bm{x}}$ with constant $L_{1}$. Denote ${\bm{f}}(\cdot,\cdot)$ the consistency function of the PF ODE. Assume the ODE solver $\Phi_{ODE}$ has a local error bound of $O((t_{n+1}-t_{n}))^{p+1}$ with $p\geq 1$. Then, if we have $\mathcal{L}_{CD}^{N}({\bm{\theta}},\Phi_{ODE})=0$, then
$\sup\limits_{n,{\bm{x}}}||{\bm{f}}_{\theta}({\bm{x}}_{t_{n}},t_{n})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n})||_{2}=O((\Delta t)^{p})$
Proof of lemma 3 in (Song et al. 2023).
From $\mathcal{L}_{CD}^{N}(\mathbf{\theta},\Phi_{ODE})=0$, derive
${\bm{f}}_{\theta}(\mathbf{x_{t_{n+1}}},t_{n+1})\equiv{\bm{f}}_{\theta}({\hat{{\bm{x}}}_{t_{n}}^{\Phi_{ODE}}},t_{n}).$
Denote
$\mathbf{e_{n+1}}$ $=\mathbf{f_{\theta}}(\mathbf{x_{t_{n+1}}},t_{n+1})-\mathbf{f}(\mathbf{x_{t_{n+1}}},t_{n+1})$
$=\mathbf{f_{\theta}}(\mathbf{\hat{x}_{t_{n}}^{\Phi_{ODE}}},t_{n})-\mathbf{f_{\theta}}(\mathbf{x_{t_{n}}},t_{n})+\mathbf{f_{\theta}}(\mathbf{x_{t_{n}}},t_{n})-\mathbf{f}(\mathbf{x_{t_{n+1}}},t_{n+1})$
$=\mathbf{f_{\theta}}(\mathbf{\hat{x}_{t_{n}}^{\Phi_{ODE}}},t_{n})-\mathbf{f_{\theta}}(\mathbf{x_{t_{n}}},t_{n})+\mathbf{f_{\theta}}(\mathbf{x_{t_{n}}},t_{n})-\mathbf{f}(\mathbf{x_{t_{n}}},t_{n})$
$=\mathbf{f_{\theta}}(\mathbf{\hat{x}_{t_{n}}^{\Phi_{ODE}}},t_{n})-\mathbf{f_{\theta}}(\mathbf{x_{t_{n}}},t_{n})+\mathbf{e_{n}}.$
Recall that $\mathbf{f_{\theta}(\cdot,\cdot)}$ has Lipschitz constant $L_{1}$,
$||e_{n+1}||_{2}$ $\leq||e_{n}||_{2}+L_{1}||\mathbf{\hat{x}_{t_{n}}^{\Phi_{ODE}}}-\mathbf{x_{t_{n}}}||_{2}$
$=||e_{n}||_{2}+L_{1}\cdot O((t_{n+1}-t_{n})^{p+1})$
$=||e_{n}||_{2}+O((t_{n+1}-t_{n})^{p+1}).$
With the boundary condition where
$\mathbf{e_{1}}=\mathbf{f_{\theta}}(\mathbf{x_{t_{1}}},t_{1})-\mathbf{f}(\mathbf{x_{t_{1}}},t_{1})=\mathbf{x_{t_{1}}}-\mathbf{x_{t_{1}}}=0,$
then
$||e_{n}||$ $\leq||e_{1}||+\sum\limits_{i=1}^{i=n-1}O((t_{i+1}-t_{i})^{p+1})$
$\leq O(({\Delta t})^{p})(T-\tau)=O(({\Delta t})^{p}).$
With this, the proof for lemma 3 is completed. ∎
Then we provide the proof for convergence of using SDE solvers in CD.
Proof of Theorem 2.
Consider the reverse SDE,
SDE:
[…]
SDE:
[…]
Following (Lu et al. 2022b), we reparameterize the SDE with $\lambda_{t}=\log\frac{\alpha_{t}}{\sigma_{t}}$, $g^{2}_{t}:=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{dlog}\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}=-2\sigma^{2}_{t}\frac{\mathrm{d}\lambda_{t}}{\mathrm{d}t}$ and $g_{t}=\sigma_{t}\sqrt{-2\frac{\mathrm{d}\lambda_{t}}{\mathrm{d}t}}$, $\mathrm{d}\mathbf{w_{\lambda}}:=\sqrt{-\frac{\mathrm{d}\lambda_{t}}{\mathrm{d}t}}\mathrm{d}\bar{\mathbf{w}}_{t}$.
Derive the integration of Equation 8 from ${\bm{x}}_{t_{n}}$ to ${\bm{x}}_{t_{m}}$, which is
$\mathbf{x_{t_{m}}}=\underbrace{\frac{\alpha_{t_{m}}}{\alpha_{t_{n}}}\mathbf{x_{t_{n}}}}_{(1)}-\underbrace{2\alpha_{t_{m}}\int_{\lambda_{t_{n}}}^{\lambda_{t_{m}}}e^{-\lambda}\mathbf{\hat{\epsilon}_{\theta}}(\mathbf{\hat{x}_{\lambda}},\lambda)\mathrm{d}\lambda}_{(2)}+\underbrace{\sigma_{t_{n}}\sqrt{e^{2(\lambda_{t_{m}}-\lambda_{t_{n}})}-1}{\bm{z}}_{t_{n}}}_{(3)},$ (9)
where ${\bm{z}}_{t_{n}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The first-order DPM SDE solver takes the form of
$\mathbf{\hat{x}^{\Phi_{SDE}}_{t_{m}}}=\underbrace{\frac{\alpha_{t_{m}}}{\alpha_{t_{n}}}\mathbf{x_{t_{n}}}}_{(1)}-\underbrace{2\sigma_{t_{m}}(e^{\lambda_{t_{m}}-\lambda_{t_{n}}}-1)\mathbf{\epsilon_{\theta}}(\mathbf{x_{t_{n}}},t_{n})}_{(2)}+\underbrace{\sigma_{t_{n}}\sqrt{e^{2(\lambda_{t_{m}}-\lambda_{t_{n}}})-1}\mathbf{z_{t_{m}}}}_{(3)}.$ (10)
Thus we can bound the local error between Equation 10 and Equation 9. Since the (1) term can be calculated exactly, the total errors $\Delta$ between Equation 10 and Equation 9 can be divided into the errors $\Delta_{2}$ in (2) term and the errors $\Delta_{3}$ in (3) term
$||{\bm{x}}_{t_{m}}-{\hat{{\bm{x}}}^{\Phi_{SDE}}_{t_{m}}}||_{2}=||\Delta||_{2}=||\Delta_{2}+\Delta_{3}||_{2}\leq||\Delta_{2}||_{2}+||\Delta_{3}||_{2}.$ (11)
Denote $h_{t_{n}}:=\lambda_{t_{m}}-\lambda_{t_{n}}$ and $\Delta_{h}:=\max\limits_{n\in[1,N]}h_{t_{n}}$. Consider the $||\Delta_{2}||$ term, we can estimate the error $\Delta_{2}$ by the Peano remainder term of the Taylor expansion,
$\int_{\lambda_{t_{n}}}^{\lambda_{t_{m}}}e^{-\lambda}\mathbf{\hat{\epsilon}_{\theta}}(\mathbf{\hat{x}_{\lambda}},\lambda)\mathrm{d}\lambda=(e^{-\lambda_{t_{n}}}-e^{-\lambda_{t_{m}}})\mathbf{\epsilon_{\theta}}(\mathbf{x_{t_{n}}},t_{n})+O(({h_{t_{n}}})^{2}).$ (12)
With this, $||\Delta_{2}||_{2}$ is bounded by $O((\Delta_{h})^{2})$.
Consider the $||\Delta_{3}||$ term, we can rewrite
$\Delta_{3}$ $=\sigma_{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}\mathbf{z_{t_{n}}}-\sigma_{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}\mathbf{z_{t_{m}}}$ (13)
$=\frac{\sigma_{t_{n}}}{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}t_{n}\mathbf{z_{t_{n}}}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}t_{m}\mathbf{z_{t_{m}}}.$ (14)
Since $\lim\limits_{h\rightarrow 0}\frac{\sqrt{e^{2h-1}}}{h}=1$, we can derive that
$\Big{|}\frac{\sigma_{t_{n}}}{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}\Big{|}=\Big{|}\frac{\sigma_{t_{n}}}{t_{n}t_{m}}\underbrace{\sqrt{e^{2h_{t_{n}}}-1}}_{O({h_{t_{n}}})}\underbrace{(t_{n}-t_{m})}_{O({h_{t_{n}}})}\Big{|},$ (15)
which is bounded by $O((\Delta_{h})^{2})$.
Thus, we can derive the bound of the $||\Delta_{3}||$ term,
$|\Delta_{3}|$ $=\Big{|}\frac{\sigma_{t_{n}}}{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}t_{n}\mathbf{z_{t_{n}}}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}t_{m}\mathbf{z_{t_{m}}}\Big{|}$ (16)
$=\Big{|}\frac{\sigma_{t_{n}}}{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}t_{n}\mathbf{z_{t_{n}}}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}t_{n}\mathbf{z_{t_{n}}}+\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}t_{n}\mathbf{z_{t_{n}}}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}t_{m}\mathbf{z_{t_{m}}}\Big{|}$ (17)
$\leq\Big{|}\frac{\sigma_{t_{n}}}{t_{n}}\sqrt{e^{2h_{t_{n}}}-1}-\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}\Big{|}\cdot\Big{|}t_{n}\mathbf{z_{t_{n}}}\Big{|}+\Big{|}\frac{\sigma_{t_{n}}}{t_{m}}\sqrt{e^{2h_{t_{n}}}-1}\Big{|}\cdot\Big{|}t_{n}\mathbf{z_{t_{n}}}-t_{m}\mathbf{z_{t_{m}}}\Big{|}.$ (18)
Notice that $t{\bm{z}}$ is a Wiener process, therefore, we have $|t_{n}\mathbf{z_{t_{n}}}-t_{m}\mathbf{z_{t_{m}}}\Big{|}$ bounded by $|O(h_{t_{n}})|$.
Thus, we have
[…]
$\leq O((h_{t_{n}})^{2}).$ (20)
With this, we prove that both $||\Delta_{2}||_{2}$ and $||\Delta_{3}||_{2}$ are bounded by $O((\Delta_{h})^{2})$. Thus we can verify the bound $O((\Delta_{h})^{2})$ of the total local errors $||\Delta||_{2}$ in Equation 11.
By lemma 3, the SDE solver has a local error bound $O((\Delta_{h})^{2})$, and $O((\Delta_{h})^{2})=O((\Delta_{t})^{2})$, we have $\sup\limits_{n,{\bm{x}}}||{\bm{f}}_{\theta}({\bm{x}}_{t_{n}},t_{n})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n})||_{2}=O((\Delta t)^{p})$, p=1, which completes the proof.
2 Multi-step Sampling of Stochastic Consistency Distillation Model